# Analysis of neutrino oscillation parameters in the light on quantum entanglement

### v1

@@ Abstract
/ Numerous neutrino experiments have confirmed the phenomenon of neutrino oscillation, providing direct evidence of the quantum mechanical nature of neutrinos. In this work, we investigate the entanglement properties of neutrino flavor states within the framework of three-flavor neutrino oscillation using two major entanglement measures: entanglement of formation (EOF) and concurrence, utilizing the DUNE experimental setup. Our findings indicate that the maximally entangled state appears between $\nu_{\mu}$ and $\nu_{\tau}$ whereas, $\nu_{e}$ behaves as a nearly separable state. To further explore the nature of bipartite entanglement, we introduce the concept of the monogamy of entanglement, which allows us to investigate the distinction between genuine three-flavor entanglement and bipartite entanglement. Our analysis confirms that the three-flavor neutrino system forms a bipartite entanglement structure, adhering to the Coffman-Kundu-Wootters (CKW) inequality. Additionally, we implement a minimization procedure to find the best-fit values of the oscillation parameters that correspond to the concurrence minima at the two specific energy points where the concurrence reaches its lowest values. Using these best-fit values, we probe three fundamental unknowns in neutrino oscillation: CP violation sensitivity, neutrino mass hierarchy, and the octant issue of $\theta_{23}$, across two distinct energy points. Our results manifest that while the best-fit values obtained through concurrence minimization show slightly reduced sensitivity to CP violation compared to current best-fit values, they exhibit greater sensitivity to the mass hierarchy. Furthermore, the study reveals a maximal mixing angle for the atmospheric sector.
@@ 1 Introduction
Given that both neutrino oscillation and quantum entanglement originate from the quantum superposition principle, investigation of potential connections between these phenomena presents a compelling research avenue Blasone:2007vw ; Blasone:2010ta ; Blasone:2013zaa ; Blasone:2014jea ; Blasone:2015lya ; Banerjee:2015mha . Extensive theoretical work has laid a foundational understanding of entangled states in neutrinos, particularly within the flavor basis Ettefaghi:2023zsh ; Blasone:2022joq ; Li:2022mus ; Patwardhan:2021rej ; Ettefaghi:2020otb . Studies have demonstrated the existence of entangled states in both bipartite and tripartite configurations of neutrino flavor states from a theoretical perspective Cervia:2019res ; Mohammed:2019svq ; Floerchinger:2018mla ; Akhmedov:2010ua ; Kayser:2010bj . Notably, concurrence is directly tied to the probabilities of the appearance and disappearance of the three neutrino flavor states Jha:2020qea ; KumarJha:2020pke during oscillation, whereas EOF represents a logarithmic function of these probabilities Konwar:2024nrd . Despite these theoretical advancements, experimentally probing the existence of entangled neutrino states and applying these concepts to address unresolved questions in neutrino physics remain formidable challenges. Several theoretical proposals have been advanced to address the unresolved questions in neutrino physics Blasone:2023qqf ; Blasone:2007vw ; Quinta:2022sgq . These approaches aim to explain key issues such as the nature of neutrino mass Dixit:2018kev ; Naikoo:2017fos , the mechanism behind neutrino oscillationsBlasone:2014cub ; Blasone:2007wp ; Siwach:2022xhx , and the potential for CP violation in the lepton sector that also include the extensions to the Standard Model in presence of non-standard neutrino interactions Konwar:2024nrd ; Dixit:2020ize . While significant progress has been made Blasone:2023qqf ; Boyanovsky:2011xq ; Ahluwalia:2011ea ; Kayser:2010bj ; Li:2021epj , a comprehensive phenomenological analysis that incorporates these aspects continues to be an open area for further exploration.
@@
In this work, we demonstrate the formation of maximally entangled states within the experimental setup of Deep Underground Neutrino Experiment (DUNE) and extend this to test the entanglement monogamy relation, establishing quantum correlations between flavor states and entanglement shareability in DUNE. Utilizing the entanglement measures EOF and concurrence, we reveal that in the DUNE setup, $\nu_{\mu}$ and $\nu_{\tau}$ form a maximally mixed state that satisfies the monogamy inequality for the three-flavor neutrino oscillation. We employ the Coffman-Kundu-Wootters (CKW) inequality to verify the monogamy relation in the entangled state, which indicates the non-shareability of entanglement between the $\nu_{\mu}$ and $\nu_{\tau}$ flavor states. With this foundational understanding of entangled flavor states in DUNE, we proceed to apply the concept of concurrence minimization to determine a set of best-fit values for the neutrino oscillation parameters in the context of quantum information. These results prompt further investigation into the experiment’s sensitivity for addressing three major questions in neutrino physics: the CP violation phase, mass hierarchy, and octant degeneracy in atmospheric mixing angle.
@@ 2 Theoretical Framework
Consider an ultra-relativistic neutrino (left-handed) with flavor $\alpha$ with $\alpha=e,\mu,\tau$ and momentum $\vec{p}$. The current theoretical framework accounts for two distinct eigenstates of neutrinos: the mass eigenstate, which is the propagating state, and the flavor eigenstate, which is the interacting state. In neutrino oscillations, the flavor state is expressed as a linear superposition of three mass eigenstates, reflecting the dual bases of neutrinos: the mass basis: |$\ket{\nu_{k}},(k=1,2,3)$ and flavor basis |$\ket{\nu_{\alpha}},(\alpha=e,\mu.\tau)$. These two bases are connected through a unitary transformation |$\ket{\nu_{\alpha}}=U\ket{\nu_{k}}$, where $U$ is the neutrino mixing matrix. We work in a basis in which charge leptons are already diagonal, the light neutrino mass matrix is diagonalised by a unitary mixing matrix known as Pontecorvo, Maki, Nakagawa, Sakata (PMNS) mixing matrix, $U_{\rm PMNS}\equiv U$ Pontecorvo:1957qd ; 10.1143/PTP.28.870 ; Pontecorvo:1967fh ; Patra:2023ltl
@@
where, $c_{ij}=\cos{\theta_{ij}}$ and $s_{ij}=\sin{\theta_{ij}}$ ($i,j=1,2,3$). Neutrino oscillations can be significantly altered in the presence of Earth’s matter. This modification arises due to the coherent forward scattering of neutrinos with particles in the medium, which introduces a potential term, $V_{CC}$, given by $V_{CC}=\sqrt{2}G_{F}N_{e}$ PhysRevD.17.2369 ; Mikheyev:1985zog ; Bethe:1986ej ; Smirnov:2004zv , where $G_{F}$ is the Fermi coupling constant and $N_{e}$ represents the electron number density. To analyze three-flavor neutrino oscillations in matter, we employ the density matrix formalism and use the $\alpha-s_{13}$ approximation Akhmedov_2004 ; Asano:2011nj to construct the qubit state of the neutrino. The density matrix formalism provides a robust framework for analyzing quantum systems, capturing both the appearance and disappearance probabilities of neutrinos. The effective Hamiltonian in the flavor basis is then expressed as,
$\mathcal{H}=\frac{\Delta m_{31}^{2}}{2E}\big{[}U\hskip 0.56917pt\mbox{diag}(0,\alpha,1)\hskip 0.56917ptU^{\dagger}+\mbox{diag}(\hat{A},0,0)\big{]},$ (1)
where $\alpha=\Delta m_{21}^{2}/\Delta m_{31}^{2}$, $\hat{A}=A/\Delta m^{2}_{31}$ with $A=2V_{CC}$. The simplified version of Eq. 1 can be written as
@@
|$\ket{\nu_{\alpha}(t)}=\sum_{\beta=e,\mu,\tau}\bigg{(}U_{\beta 1}U^{*}_{\alpha 1}e^{-iE_{1}t}+U_{\beta 2}U^{*}_{\alpha 2}e^{-iE_{2}t}+U_{\beta 3}U^{*}_{\alpha 3}e^{-iE_{3}t}\bigg{)}\ket{\nu_{\beta}},$ (3)
where $U_{\alpha i}$’s are the component of the matter-modified mixing matrix $U_{M}$. This leads to the construction of the time-evolved flavor state taking $\nu_{\mu}$ as initial flavor state,
|$\ket{\nu_{\mu}(t)}=\tilde{U}_{\mu e}\ket{\nu_{e}}+\tilde{U}_{\mu\mu}\ket{\nu_{\mu}}+\tilde{U}_{\mu\tau}\ket{\nu_{\tau}},$ (4)
@@
$\rho_{\mu}(t)=\begin{pmatrix}0&0&0&0&0&0&0&0\\ 0&~{}|\tilde{U}_{\mu\tau}|^{2}~{}&~{}\tilde{U}_{\mu\tau}\tilde{U}^{*}_{\mu\mu}~{}&~{}0~{}&~{}\tilde{U}_{\mu\tau}\tilde{U}^{*}_{\mu e}~{}&~{}0~{}&~{}0~{}&0\\ 0&~{}\tilde{U}_{\mu\mu}\tilde{U}^{*}_{\mu\tau}~{}&~{}|\tilde{U}_{\mu\mu}|^{2}~{}&~{}0~{}&~{}\tilde{U}_{\mu\mu}\tilde{U}^{*}_{\mu e}~{}&~{}0&0&0\\ 0&0&0&0&0&0&0&0\\ 0&\tilde{U}_{\mu e}\tilde{U}^{*}_{\mu\tau}&\tilde{U}_{\mu e}\tilde{U}^{*}_{\mu\mu}&0&|\tilde{U}_{\mu e}|^{2}&0&0&0\\ 0&0&0&0&0&0&0&0\\ 0&0&0&0&0&0&0&0\\ 0&0&0&0&0&0&0&0\\ \end{pmatrix}.\$ (5)
Clearly from the above expression $|\tilde{U}_{\mu\mu}|^{2}=P_{\mu\mu}$, $|\tilde{U}_{\mu e}|^{2}=P_{\mu e}$ and $|\tilde{U}_{\mu\tau}|^{2}=P_{\mu\tau}$ with the normalization condition $|\tilde{U}_{\mu\mu}|^{2}+|\tilde{U}_{\mu e}|^{2}+|\tilde{U}_{\mu\tau}|^{2}=1$. The classification of entanglement in a three-qubit system can be categorized as follows:
@@
Entanglement of formation (EOF): The entanglement of formation is the most basic measurement of entanglement. Given a density matrix $\rho$ of a pair of quantum systems A and B, consider all possible pure state decompositions of $\rho$, i.e., all ensembles of states |$\ket{\psi_{i}}$ with probabilities $p_{i}$, such that
$\rho=\sum p_{i}\ket{\psi_{i}}\bra{\psi_{i}}.$ | . (6)
@@
Here, $\rho_{A}$ is the partial trace of |$\ket{\psi}\bra{\psi}$ | over subsystem B, and $\rho_{B}$ is defined similarly PhysRevLett.70.1895 ; PhysRevA.54.3824 . From 7, the term $\mbox{Tr}(\rho_{A}\log_{2}\rho_{A})$ effectively suggests the entropy of the pure system $A$. Consequently, the expression for the relative entropy for a bipartite system is $-\sum_{i}k_{i}^{2}\log_{2}k_{i}^{2}$ PhysRevLett.80.2245 , where $k_{i}$ is the Schmidt coefficient and satisfies the condition $\sum_{i}k_{i}^{2}=1$. Analogically, with the neutrino system the Schmidt coefficient $k^{2}_{i}$ leads to the appearance and disappearance probabilities $P_{a}$ and $P_{d}$ adhering the property ($P_{d}+\sum P_{a})=1$. This observation leads to the same conclusion about the conservation of probabilities. For a tripartite system, the EOF is given by PhysRevA.99.042305 ; PhysRevA.101.032301 ,
$\begin{split}{\rm EOF}&=(S_{A}+S_{B}+S_{C}),\\ \end{split}$ (8)
@@
$C^{\alpha}=\big{[}3-{\rm Tr}(\rho_{\alpha})^{2}+{\rm Tr}(\rho_{\beta})^{2}+{\rm Tr}(\rho_{\gamma})^{2}\big{]}^{1/2}.$ (10)
For the three-flavor neutrino system Eq. 10 modifies to the expression,

### v2

@@ Abstract
NNumerous neutrino experiments have confirmed the phenomenon of neutrino oscillation, providing direct evidence of the quantum mechanical nature of neutrinos. In this work, we investigate the entanglement properties of neutrino flavor states within the framework of three-flavor neutrino oscillation using two major entanglement measures: entanglement of formation (EOF) and concurrence, utilizing the DUNE experimental setup. Our findings indicate that the maximally entangled state appears between $\nu_{\mu}$ and $\nu_{\tau}$ whereas, $\nu_{e}$ behaves as a nearly separable state. To further explore the nature of bipartite entanglement, we introduce the concept of the monogamy of entanglement, which allows us to investigate the distinction between genuine three-flavor entanglement and bipartite entanglement. Our analysis confirms that the three-flavor neutrino system forms a bipartite entanglement structure, adhering to the Coffman-Kundu-Wootters (CKW) inequality. Additionally, we implement a minimization procedure to find the best-fit values of the oscillation parameters that correspond to the concurrence minima at the two specific energy points where the concurrence reaches its lowest values. Using these best-fit values, we probe three fundamental unknowns in neutrino oscillation: CP violation sensitivity, neutrino mass hierarchy, and the octant issue of $\theta_{23}$, across two distinct energy points. Our results manifest that while the best-fit values obtained through concurrence minimization show slightly reduced sensitivity to CP violation compared to current best-fit values, they exhibit greater sensitivity to the mass hierarchy. Furthermore, the study reveals a maximal mixing angle for the atmospheric sector.
@@ 1 Introduction
Given that both neutrino oscillation and quantum entanglement originate from the quantum superposition principle, investigation of potential connections between these phenomena presents a compelling research avenue Blasone:2007vw; Blasone:2010ta; Blasone:2013zaa; Blasone:2014jea; Blasone:2015lya; Banerjee:2015mha. Extensive theoretical work has laid a foundational understanding of entangled states in neutrinos, particularly within the flavor basis Ettefaghi:2023zsh; Blasone:2022joq; Li:2022mus; Ettefaghi:2020otb. In the neutrino sector, it is a well-established concept that flavor eigenstates, corresponding to the interaction eigenstates, are quantum superpositions of mass eigenstates, also known as propagation eigenstates. This relationship is particularly evident at the moment of neutrino production, where the neutrino exists in a pure quantum state. During propagation, however, the interaction between the flavor and mass bases results in a dynamic interchange, wherein the mass eigenstates acquire flavor compositions and vice versa. Consequently, the flavor eigenstates evolve into superpositions of different flavor components Smirnov:2003da. Notably, this superposition is not a simple quantum superposition but may instead give rise to an entangled superposition state Blasone:2007vw; Blasone:2007wp. The fundamental distinction between a simple quantum superposition and an entangled superposition lies in the fact that the latter cannot be factorized into a direct product of two independent component states.
Studies have demonstrated the existence of entangled states in both bipartite and tripartite configurations of neutrino flavor states from a theoretical perspective Bittencourt:2023asd; Mohammed:2019svq; Floerchinger:2018mla; Akhmedov:2010ua; Kayser:2010bj. This work aims to investigate the possibility of the existence of such entangled superposition states in the context of neutrino oscillations, providing new insights into their quantum behavior and underlying structure in the experimental aspect. Since the entanglement arises through the dynamical changes of the flavor states, it is reasonable to verify the amount of entanglement in terms of two entanglement-measures such as EOF and concurrence. This is because EOF represents a logarithmic function of these probabilities Konwar:2024nrd whereas, concurrence is directly tied to the probabilities of the appearance and disappearance of the three neutrino flavor states Jha:2020qea; KumarJha:2020pke during oscillation.
@@
Despite these theoretical advancements, experimentally probing the existence of entangled neutrino states and applying these concepts to address unresolved questions in neutrino physics remain formidable challenges. Several theoretical proposals have been advanced to address the unresolved questions in neutrino physics Blasone:2023qqf; Blasone:2007vw; Quinta:2022sgq. These approaches aim to explain key issues such as the nature of neutrino mass Dixit:2018kev; Naikoo:2017fos, the mechanism behind neutrino oscillations Blasone:2014cub; Blasone:2007wp, and the potential for CP violation in the lepton sector that also include the extensions to the Standard Model in presence of non-standard neutrino interactions Konwar:2024nrd; Dixit:2020ize. While significant progress has been made Blasone:2023qqf; Boyanovsky:2011xq; Ahluwalia:2011ea; Kayser:2010bj; Li:2021epj, a comprehensive phenomenological analysis that incorporates these aspects continues to be an open area for further exploration.
@@
In this work, we demonstrate the formation of maximally entangled states within the experimental setup of Deep Underground Neutrino Experiment (DUNE) and extend this to test the entanglement monogamy relation, establishing quantum correlations between flavor states and entanglement shareability in DUNE. Utilizing the entanglement measures EOF and concurrence, we reveal that in the DUNE setup, $\nu_{\mu}$ and $\nu_{\tau}$ form a maximally mixed state that satisfies the monogamy inequality for the three-flavor neutrino oscillation. We employ the Coffman-Kundu-Wootters (CKW) inequality to verify the monogamy relation in the entangled state, which indicates the non-shareability of entanglement between the $\nu_{\mu}$ and $\nu_{\tau}$ flavor states. With this foundational understanding of entangled flavor states in DUNE, we proceed to apply the concept of concurrence minimization to determine a set of best-fit values for the neutrino oscillation parameters in the context of quantum information. These results prompt further investigation into the experiment’s sensitivity for addressing three major questions in neutrino physics: the CP violation, mass hierarchy, and octant degeneracy in atmospheric mixing angle.
@@ 2 Theoretical Framework
[…]
@@
where, $c_{ij}=\cos{\theta_{ij}}$ and $s_{ij}=\sin{\theta_{ij}}$ ($i,j=1,2,3$). Neutrino oscillations can be significantly altered in the presence of Earth’s matter. This modification arises due to the coherent forward scattering of neutrinos with particles in the medium, which introduces a potential term, $V_{CC}$, given by $V_{CC}=\sqrt{2}G_{F}N_{e}$ PhysRevD.17.2369; Mikheyev:1985zog; Bethe:1986ej; Smirnov:2004zv. To analyze three-flavor neutrino oscillations in matter, we employ the density matrix formalism and use the $\alpha-s_{13}$ approximation Akhmedov_2004; Asano:2011nj to construct the qubit state of the neutrino. The density matrix formalism provides a robust framework for analyzing quantum systems, capturing both the appearance and disappearance probabilities of neutrinos. The effective Hamiltonian in the flavor basis is then expressed as,
$\mathcal{H}=\frac{\Delta m_{31}^{2}}{2E}\left[U\hskip 0.56917pt\mbox{diag}(0,\alpha,1)\hskip 0.56917ptU^{\dagger}+\mbox{diag}(\hat{A},0,0)\right],$ (1)
where $\alpha=\Delta m_{21}^{2}/\Delta m_{31}^{2}$, $\hat{A}=A/\Delta m^{2}_{31}$ with $A=2V_{CC}$. It is to be mentioned here that in the three-flavor neutrino oscillation framework, only Charged-Current (CC) interactions are included, as they distinguish neutrino flavors through coupling with their corresponding charged leptons via $W^{\pm}$ exchange. Neutral-Current (NC) interactions, mediated by the $Z$ boson, are flavor-independent and contribute an equal potential to all flavors, introducing only an overall phase that does not affect oscillation probabilities. Hence, the NC term is omitted in the effective Hamiltonian without loss of generality. The simplified version of Eq.(1) can be written as
$\mathcal{H}=\frac{\Delta m_{31}^{2}}{2E}R_{23}\hskip 0.56917ptU_{\delta}\hskip 0.56917ptM\hskip 0.56917ptU_{\delta}^{\dagger}\hskip 0.56917ptR_{23}^{T},$ (2)
where, $U_{\delta}=\mbox{diag}(1,1,e^{i\delta_{CP}})$ and $M=[R_{13}\hskip 0.56917ptR_{12}\hskip 0.56917pt\mbox{diag}(0,\alpha,1)\hskip 0.56917ptR_{12}^{T}\hskip 0.56917ptR_{13}^{T}+\mbox{diag}(\hat{A},0,0)]$. Diagonalization of this matrix yields the eigenvalues $E_{1}$, $E_{2}$ and $E_{3}$ and the corresponding eigenvectors $v_{1}$, $v_{2}$ and $v_{3}$. By stacking these eigenvectors, the resulting mixing matrix is given by $W=(v_{1}\quad v_{2}\quad v_{3})$. The modified mixing matrix in matter is then expressed as $\mathbb{U}=R_{23}\hskip 0.56917ptU_{\delta}\hskip 0.56917ptW$. The time-evolved flavor state is given by Konwar:2024nrd
@@
where, $\mathbb{U}_{\alpha i}$’s are the component of the matter-modified mixing matrix $\mathbb{U}$ and $\tilde{\mathbb{U}}_{\alpha\beta}=\mathbb{U}_{\beta i}\mathbb{U}_{\alpha i}^{*}e^{-iE_{i}t}$. This leads to the construction of the time-evolved flavor state taking $\nu_{\mu}$ as initial flavor state,
@@
[…]
This perspective motivates a further investigation into neutrino entanglement, particularly in the context of the DUNE experiment. For DUNE set up, one prepares muon neutrino beam $\nu_{\mu}$ which propagates through the medium and finally, evolves into different flavor modes as $\nu_{e},\nu_{\mu},\nu_{\tau}$. In this scenario, the density matrix for the state $|\rho_{\mu}(t)=\ket{\nu_{\mu}(t)}\bra{\nu_{\mu}(t)}$, is given by Konwar:2024nrd; Li:2021epj,
$\rho_{\mu}(t)=\begin{pmatrix}0&0&0&0&0&0&0&0\\ 0&~|\widetilde{\mathbb{U}}_{\mu\tau}|^{2}~&~\widetilde{\mathbb{U}}_{\mu\tau}\widetilde{\mathbb{U}}^{*}_{\mu\mu}~&~0~&~\widetilde{\mathbb{U}}_{\mu\tau}\widetilde{\mathbb{U}}^{*}_{\mu e}~&~0~&~0~&0\\ 0&~\widetilde{\mathbb{U}}_{\mu\mu}\widetilde{\mathbb{U}}^{*}_{\mu\tau}~&~|\widetilde{\mathbb{U}}_{\mu\mu}|^{2}~&~0~&~\widetilde{\mathbb{U}}_{\mu\mu}\widetilde{\mathbb{U}}^{*}_{\mu e}~&~0&0&0\\ 0&0&0&0&0&0&0&0\\ 0&\widetilde{\mathbb{U}}_{\mu e}\widetilde{\mathbb{U}}^{*}_{\mu\tau}&\widetilde{\mathbb{U}}_{\mu e}\widetilde{\mathbb{U}}^{*}_{\mu\mu}&0&|\widetilde{\mathbb{U}}_{\mu e}|^{2}&0&0&0\\ 0&0&0&0&0&0&0&0\\ 0&0&0&0&0&0&0&0\\ 0&0&0&0&0&0&0&0\\ \end{pmatrix}$.\\ (5)
Clearly from the above expression $|\widetilde{\mathbb{U}}_{\mu\mu}|^{2}=P_{\mu\mu}$, $|\widetilde{\mathbb{U}}_{\mu e}|^{2}=P_{\mu e}$ and $|\widetilde{\mathbb{U}}_{\mu\tau}|^{2}=P_{\mu\tau}$ with the normalization condition $|\widetilde{\mathbb{U}}_{\mu\mu}|^{2}+|\widetilde{\mathbb{U}}_{\mu e}|^{2}+|\widetilde{\mathbb{U}}_{\mu\tau}|^{2}=1$. The classification of entanglement in a three-qubit system can be categorized as follows:
@@
[…]
$|.\rho=\sum p_{i}\ket{\psi_{i}}\bra{\psi_{i}}$. (6)
@@
Here, $\rho_{A}$ is the partial trace of $|\ket{\psi}\bra{\psi}$ over subsystem B, and $\rho_{B}$ is defined similarly PhysRevLett.70.1895; PhysRevA.54.3824. From (7), the term $\mbox{Tr}(\rho_{A}\log_{2}\rho_{A})$ effectively suggests the entropy of the pure system $A$. Consequently, the expression for the relative entropy for a bipartite system is $\sum_{i}k_{i}^{2}\log_{2}k_{i}^{2}$ PhysRevLett.80.2245, where $k_{i}$ is the Schmidt coefficient and satisfies the condition $\sum_{i}k_{i}^{2}=1$. Analogically, with the neutrino system the Schmidt coefficient $k^{2}_{i}$ leads to the appearance and disappearance probabilities $P_{a}$ and $P_{d}$ adhering the property ($P_{d}+\sum P_{a})=1$. This observation leads to the same conclusion about the conservation of probabilities. For a tripartite system, the EOF is given by PhysRevA.99.042305; PhysRevA.101.032301,
[…]
@@
$C^{\alpha}=\big[3-{\rm Tr}(\rho_{\alpha})^{2}-{\rm Tr}(\rho_{\beta})^{2}-{\rm Tr}(\rho_{\gamma})^{2}\big]^{1/2}.$ (10)
For the three-flavor neutrino system Eq. (10) modifies to the expression,